How do I solve:
1-4(2X+3)=5(X-2)-3(2X-1)
Thank you

Answer by donnieb82(2)

Simplifying
1 + -4(2x + 3) = 5(x + -2) + -3(2x + -1)
Reorder the terms:
1 + -4(3 + 2x) = 5(x + -2) + -3(2x + -1)
1 + (3 * -4 + 2x * -4) = 5(x + -2) + -3(2x + -1)
1 + (-12 + -8x) = 5(x + -2) + -3(2x + -1)
Combine like terms: 1 + -12 = -11
-11 + -8x = 5(x + -2) + -3(2x + -1)
Reorder the terms:
-11 + -8x = 5(-2 + x) + -3(2x + -1)
-11 + -8x = (-2 * 5 + x * 5) + -3(2x + -1)
-11 + -8x = (-10 + 5x) + -3(2x + -1)
Reorder the terms:
-11 + -8x = -10 + 5x + -3(-1 + 2x)
-11 + -8x = -10 + 5x + (-1 * -3 + 2x * -3)
-11 + -8x = -10 + 5x + (3 + -6x)
Reorder the terms:
-11 + -8x = -10 + 3 + 5x + -6x
Combine like terms: -10 + 3 = -7
-11 + -8x = -7 + 5x + -6x
Combine like terms: 5x + -6x = -1x
-11 + -8x = -7 + -1x
Solving
-11 + -8x = -7 + -1x
Solving for variable 'x'.
Move all terms containing x to the left, all other terms to the right.
Add 'x' to each side of the equation.
-11 + -8x + x = -7 + -1x + x
Combine like terms: -8x + x = -7x
-11 + -7x = -7 + -1x + x
Combine like terms: -1x + x = 0
-11 + -7x = -7 + 0
-11 + -7x = -7
Add '11' to each side of the equation.
-11 + 11 + -7x = -7 + 11
Combine like terms: -11 + 11 = 0
0 + -7x = -7 + 11
-7x = -7 + 11
Combine like terms: -7 + 11 = 4
-7x = 4
Divide each side by '-7'.
x = -0.5714285714
Simplifying
x = -0.5714285714